Problem 30.28 What is the resistance of

Part B

A 20.0cm long piece of carbon with a 0.900mm
Answer #1

Resistivity Of carbon

p=35*10-6 ohms/m

Area

A=(0.9*10-3)2

Resistance

R=pL/A =(35*10-6)(0.2)/(0.9*10-3)2

R=8,64 ohms
Answer #2

length= 20 cm = .2 meter

area = .81 mm^2 =8.1*10^-7 m^2

resistivity =1.68